OpenStudy (anonymous):

show all work. The property restoration company PuroServ is considering switching to new dehumidifiers. Their market research, considering the cost of the new machines and their efficiency, tells them that the switch would give them a 54% chance of making a $20,000 profit, a 23% chance of breaking even, and a 23% chance of losing $5,000. How much money does PuroServ expect to make with their new purchase?

OpenStudy (anonymous):

\[20000\times .54-5000\times .23\]

OpenStudy (anonymous):

10800-1150=9650